OpenStudy (richgall____):

What is the volume of a cone with a diameter of 18 cm and a height of 10 cm? Use 3.14 for pi. Round your answer to the nearest tenth.

pooja195 (pooja195):

\[\huge~\rm~V= \pi~r^2 \frac{h }{ 3}\] plug in your numbers and solve :)

pooja195 (pooja195):

Remember to find the radius since it is not given

OpenStudy (richgall____):

radius=9 right? i remember learning back in 6th grade that it;s just half of the diameter

pooja195 (pooja195):

Correct

OpenStudy (anonymous):

remember,radius is half of its diameter yea correct

pooja195 (pooja195):

\[\huge~\rm~V= (3.14)~(9)^2 (\frac{10 }{ 3})\]

OpenStudy (richgall____):

848.230016469 im using the google calculator because i dnt have my texas instrument right now so i dont think this is right

pooja195 (pooja195):

Round it to the nearest tenth

pooja195 (pooja195):

Since that's what the problem asks for

OpenStudy (richgall____):

848.2

pooja195 (pooja195):

Well done :)

OpenStudy (richgall____):

wait so the google calculator was right?

pooja195 (pooja195):

That has it to the nearest hundredth

pooja195 (pooja195):

Not tenth so if you choose to use that becareful

OpenStudy (richgall____):

could you check it for me this work is so confusing without the TI-84

pooja195 (pooja195):

Sure :)

OpenStudy (richgall____):

Okay.

pooja195 (pooja195):

Wait no that's incorrect....google calculator doesn't use pi :-/ Don't use that this is correct: 847.8

OpenStudy (richgall____):

Thanks so much i really need to go get my calculator from my friend like right now -_-

pooja195 (pooja195):

xD thanks @MARC_ for reminding me :P

OpenStudy (anonymous):

np